# Latent Action Pretraining From Videos

### v1

@@ 6 Limitations and Conclusion
In this paper, we introduce LAPA, a scalable pretraining method for building VLAs using actionless videos. Across three benchmarks spanning both simulation and real-world robot experiments, we show that our method significantly improves transfer to downstream tasks compared to existing approaches. We also present a state-of-the-art VLA model that surpasses current models trained on 970K action-labeled trajectories. Furthermore, we demonstrate that LAPA can be applied purely on human manipulation videos, where explicit action information is absent, and the embodiment gap is substantial. We believe our work can be extended to build scalable robot foundation models.
We still face certain limitations. First, LAPA underperforms compared to action pretraining when it comes to fine-grained motion generation tasks like grasping. We believe that increasing the latent action generation space could help address this issue. Second, similar to prior VLAs, LAPA also encounters latency challenges during real-time inference. Adopting a hierarchical architecture, where a smaller head predicts actions at a higher frequency, could potentially reduce latency and improve fine-grained motion generation. Lastly, while we qualitatively demonstrate that our latent action space captures camera movements (Figure 10), we have not yet explored the application of LAPA beyond manipulation videos, such as those from self-driving cars, navigation, or landscape scenes. We leave these explorations for future work. We hope that our work can help overcome the data bottleneck in robotics and accelerate the development of generalist robot policies.
Acknowledgements
We thank Arhan Jain and Marius Memmel for helping out with the robot hardware and teleoperation setup. Also, we thank Minyoung Hwang, Jiafei Duan, Junsu Kim, and Changyeon Kim for helpful discussions and constructive feedback.
Appendix A Latent Action Quatization Model Details
@@
We show model architecture details of our latent action quantization model in Figure 13. We utilize the C-ViViT model architecture from Villegas et al. (2022) to replicate the latent action model from GENIE (Bruce et al., 2024). After latent model training, we utilize the $z_{2}$ as the latent action label for $x_{1}$. The encoder can be seen as the inverse dynamics model and the decoder can be seen as the world model.
@@ Appendix B Details on Experimental Setup
Figure 3 (C) shows examples of the real-world tabletop manipulation experimental setup. For the teleoperation of single arm robot experiments, we use the polymetis robotic stack [^8] https://github.com/facebookresearch/polymetis to collect 150 trajectories for each of the tasks. All of the tasks require multi-instruction following capabilities since there are 3 objects in the scene and the model has to condition on the task description to infer which object to interact with. Figure 14 shows samples of each task. For each task, we aim to quantify 3 distinct capabilities:
(1) We test the ability to infer the correct object from the task description between an unseen combination of seen objects during fine-tuning, (2) We test the ability to infer the correct object from totally unseen objects during fine-tuning that may or may have not been observed during pretraining. Specifically, the knocking tasks was conducted with real-world objects that were highly unlikely to have been in any of the pertaining datasets. (3) We test the ability to infer the correct object (among seen objects, unseen combinations) from a totally unseen instruction that requires semantic reasoning (e.g. Pick up a spicy object). For each evaluation criteria, 6 rollouts are performed for each models, resulting in a total of 18 rollouts for each task category. Since there are three tasks, each model is evaluated with 54 rollouts in the real-world. We provide the full list of all of the seen and unseen objects used for each rollout in Table 13, 14, 15, and the total average success rates in Table 16.
@@
Similar to single arm experiments, we collect 150 trajectories for bi-manual experiments. For the bi-manual task, one of the arms should choose one between two plates depending on the language instruction and put the chosen plate on the container. Next, the other arm should pick up a correct object from three object candidates and put the object on the plate. For evaluation, we quantify (1) seen combination of seen objects during fine-tuning, (2) unseen combinations of seen objects, (3) unseen objects, (4) unseen instructions of seen objects. 6 rollouts are performed for each sub-category, resulting in 24 rollouts for each model. We provide the full result in Table 17. For evaluation metrics, we adapt a partial success criteria for fine-grained evaluation.
Put the $<$object1$>$ on the container and $<$object2$>$ on the plate.
For bi-manual task, we give 0.25 partial score if the robot reaches to the correct plate, 0.5 if the robot puts the correct plate on the container successfully, 0.75 if the robot reaches to the correct object on the right, and 1 if the robot successfully pick the object and place the object on top of the plate.
@@ Appendix D Experimental Result Analysis
We further analyze the real-world robot results shown in Figures 5 and 6(b), focusing on how the task distribution in pretraining data impacts downstream performance. Table 3 presents the number of trajectories corresponding to each evaluation task (Knocking, Covering, and Pick & Place) across pretraining datasets (Bridgev2, Open-X, and Something Something V2 (Sthv2)), determined through lexical matching. We expect future work to use other methods of analyzing the relationship between pertaining and fine-tuning task distributions that capture semantics of the task rather than simple lexical matching. We perform this analysis to get a sense of how the task distribution in the pretraining data affects downstream task performance.
@@
Appendix E Detailed Experimental Results
E.1 Language Table
We provide the detailed results of the experiments performed on the Language Table benchmark in Table 4, 5, 6, 7, 8, 9. For all of the tables in the appendix, we bold the best result among the comparisons and underline the second best. Each value denotes the success rate (%). 50 evaluation rollouts are performed for each task category, resulting in 250 total evaluation rollouts per model for each table.
@@ E.1 Language Table
For ablation of LAPA in Language Table, we experiment with various latent action vocab and sequence sizes. As shown in Figure 16, increasing the sequence and vocab size increases the performance. However, unlike SIMPLER, we observe that the increasing the latent action vocab size is much more effective compared to increasing the latent action sequence length in terms of absolute performance. This implies that for environments that are visually simple, increasing the latent action vocabulary might be more effective compared to sequence length.
@@
E.2 SIMPLER
We provide detailed results of various models evaluated on SIMPLER environment. Table 10 shows the setting where baseline models are pretrained on Bridgev2 and then finetuned on SIMPLER rollouts (100 videos). The results show detailed results for each task (stack green to yellow block, put carrot on plate, put spoon on otowel, put eggplant in basket) and subtasks (grasping and moving).
@@ E.2 SIMPLER
Table 10: SIMPLER results of Bridgev2 Pretraining. Success, Grasping, and Moving Rates (%) in SIMPLER environment. We pretrain UniPi, Vpt, and LAPA on Bridgev2 dataset without using ground-truth action labels and ActionVLA on Bridgev2 using action labels. The main 4 tasks are: stack green to yellow block, put carrot on plate, put spoon on towel, and put eggplant in basket. Best is bolded and second best is underlined.
Success Rate Scratch UniPi Vpt LAPA ActionVLA
Stack G2Y 29.2 2.7 45.8 54.2 75.0
Carrot2Towel 29.2 2.7 37.5 45.8 58.0
Spoon2Plate 50.0 0.0 70.8 70.8 70.8
Eggplant2Bask 29.2 0.0 50.0 58.3 50.0
AVG 34.4 1.3 51.0 57.3 63.5
@@
Grasp Green Block 66.6 20.8 62.5 62.5 87.5
Grasp Carrot 45.8 33.2 54.1 58.3 75.0
Grasp Spoon 70.8 22.2 79.2 83.3 83.3
Grasp Eggplant 62.5 16.0 70.8 83.3 75.0
AVG 61.4 23.1 66.7 71.9 80.2
@@
Move Green Block 58.3 29.1 58.3 66.6 91.6
Move Carrot 45.8 48.6 66.6 70.8 91.6
Move Spoon 70.8 34.6 79.2 83.3 79.2
Move Eggplant 87.5 58.0 70.8 87.5 91.6
AVG 65.6 42.6 68.7 77.1 88.5
@@
Success Rate VPT UniPi LAPA
StackG2Y 50.0 0.0 50.0
Carrot2Towel 29.1 1.3 50.0
Spoon2Plate 37.5 1.3 50.0
Eggplant2Bask 66.6 0.0 58.3
AVG 45.8 0.7 52.1
@@
Grasp Green Block 66.6 2.7 58.3
Grasp Carrot 45.8 31.7 62.5
Grasp Spoon 70.8 21.7 75.0
Grasp Eggplant 91.6 6.8 70.8
AVG 68.7 15.7 66.7
@@
Move Green Block 62.5 2.7 62.5
Move Carrot 58.3 37.5 70.8
Move Spoon 54.1 18.1 75.0
Move Eggplant 91.6 50.3 83.3
AVG 66.6 27.1 72.9
E.3 Real-world
We provide the detailed result of real world evaluation depending on the generalization type: (1) seen objects but unseen combinations, (2) unseen objects, and (3) seen objects but unseen instructions. The results are shown in Table 12. As shown in the table, LAPA (Open-X) outperforms OpenVLA (Open-X) on all types of generalization settings. Also, LAPA (Human Videos) shows good generalization performance, especially for unseen objects. We conjecture that this is because Something Something V2 dataset interacts with much diverse objects compared to Bridgev2.
@@ E.3 Real-world
Seen Objects Unseen Seen Objects AVG
Unseen Combinations Objects Unseen Instructions
Scratch 18.0 20.3 25.4 21.2
ActionVLA (Bridge) 38.3 31.8 27.7 32.6
OpenVLA (Bridge) 35.6 34.6 22.1 30.8
LAPA (Bridge) 43.4 31.4 35.6 36.8
OpenVLA (Open-X) 46.2 42.1 43.4 43.9
LAPA (Open-X) 57.8 43.9 48.5 50.1
LAPA (Human Videos) 36.5 37.4 28.1 34.0
@@
OpenVLA LAPA OpenVLA ActionVLA LAPA Scratch LAPA
@@
Unseen Object Combinations
@@
Unseen Instructions
@@
OpenVLA LAPA OpenVLA ActionVLA LAPA Scratch LAPA
@@
Unseen Object Combinations
@@
Unseen Instructions
@@
OpenVLA LAPA OpenVLA ActionVLA LAPA Scratch LAPA
@@
Unseen Object Combinations
@@
Unseen Instructions
@@
Table 16: Summary of Single Arm Total Success Rates (%)
OpenVLA LAPA OpenVLA ActionVLA LAPA Scratch LAPA
@@
Table 17: Bi-manual Task Results (Seen and Unseen)
OpenVLA (Open-X) LAPA (Open-X)
Seen Object Combinations
purple, blue cup 0.25 0.25
gray, cupcake 0.25 0.75
gray, apple 0 0
purple, milk 0.25 0.75
gray, orange juice 0.5 0.25
purple, green cup 0.25 0.25
SUM 1.5 2.25
Unseen Object Combinations
purple, orange 0.75 0
purple, red block 0.25 0
purple, cupcake 0.25 0
gray, peach 0.25 0.5
gray, mustard 0 0
gray, spam 0.25 0.5
SUM 1.75 1
Unseen Objects
purple, pear 0.25 0.5
purple, mayonnaise 0.75 0.25
purple, yellow cup 0.5 0.5
plain white, tomato soup 0.25 0.75
plain white, wooden block 0.25 0.25
plain white, cheez it 0 0
SUM 2 2.25
Unseen Instructions
darker, fruit 0.25 0.5
darker, sweet dessert 0.5 0
darker, drink 0.25 0.5
lighter, rectangular object 0 0.25
lighter, roundest object 0 0.5
lighter, red object 0 0
SUM 1 1.75
Success Rate 26.04% 30.21%

### v2

@@ 6 Limitations and Conclusion
In this paper, we introduce Latent Action Pretraining, a scalable pretraining method for building VLAs without using ground-truth action labels. Across three benchmarks spanning both simulation and real-world robot experiments, we show that our method significantly improves transfer to downstream tasks compared to existing approaches. We also present a state-of-the-art VLA model that surpasses current models trained on 970K action-labeled trajectories. Furthermore, we demonstrate that our method can be applied purely on human manipulation videos, where explicit action information is absent, and the embodiment gap is substantial.
We still face certain limitations. First, LAPA underperforms compared to action pretraining when it comes to fine-grained motion generation tasks like grasping. We believe that increasing the latent action generation space could help address this issue. Second, similar to prior VLAs, LAPA also encounters latency challenges during real-time inference. Adopting a hierarchical architecture, where a smaller head predicts actions at a higher frequency. Lastly, while we qualitatively demonstrate that our latent action space captures camera movements (Figure 14), we have not yet explored the application of LAPA beyond manipulation videos, such as those from self-driving cars, navigation, or landscape scenes. We leave these explorations for future work.
Acknowledgments
We thank Arhan Jain and Marius Memmel for helping out with the robot hardware and teleoperation setup. Also, we thank Minyoung Hwang, Jiafei Duan, Junsu Kim, and Changyeon Kim for helpful discussions and constructive feedback. This work was partly supported by Center for Advanced Urban Systems (CAUS) of Korea Advanced Institute of Science and Technology (KAIST) funded by GS E&C (40%) and the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2024-00397966, Development of a Cybersecurity Specialized RAG-based sLLM Model for Suppressing Gen-AI Malfunctions and Construction of a Publicly Demonstration Platform, 30%; No.RS-2022-II220264, Comprehensive Video Understanding and Generation with Knowledge-based Deep Logic Neural Network, 20%; No.RS-2021-II212068, Artificial Intelligence Innovation Hub, 10%).
Appendix A Latent Action Quantization Details
@@
We show model architecture details of our latent action quantization model in Figure 8. We utilize the C-ViViT model architecture from Villegas et al. (2022) to replicate the latent action model from GENIE (Bruce et al., 2024). During the encoding process, both $x_{1}$ and $x_{2}$ are gone through a patch embedding to obtain $p_{1}$ and $p_{2}$ and gone through a spatial transformer. To convey temporal information, the representations for both outputs of spatial transformer are passed to a causal transformer (transformer with causal positional encodings) to get $e_{1}$ and $e_{2}$ continuous embeddings. We then define $d_{1}=e_{2}-e_{1}$ and discretize $d_{1}$ by finding the closest embedding from the codebook $z$ where the codebook size is determined as a hyperparameter.
$z_{1}=\arg\min_{z_{k}}|d_{1}-z_{k}|^{2}$ (1)
When obtaining $e_{1}$ and $e_{2}$, we go through a CNN network. The sequence length is designated by the kernel size, stride and padding value of a CNN network. After quantization, we apply NSVQ technique before decoding.
$\hat{d_{1}}=d_{1}+\frac{\|d_{1}-z_{1}\|}{\|v\|}v$ (2)
where $v\sim\mathcal{N}(0,1)$. For decoding, we go through the following equation:
$\hat{x_{2}}=D(Attn(sg[p_{1}],\hat{d_{1}},\hat{d_{1}})$ (3)
where stop gradient ($sg$) is applied to $p_{1}$ to avoid representation collapse and cross attention is used to attend $\hat{d_{1}}$ given $p_{1}$. Unlike the encoder, decoder $D$ only include spatial transformer. The training objective is a L2 reconstruction loss.
$L=\|x_{2}-\hat{x_{2}}\|_{2}^{2}$ (4)
After latent model training, we utilize the $z_{1}$ as the latent action label for $x_{1}$. The encoder can be seen as the inverse dynamics model and the decoder can be seen as the world model.
@@ Appendix B Details on Experimental Setup
Table 3: Pretraining and fine-tuning dataset for each environment. Cross-Env denotes cross-environment, Cross-Emb denotes cross-embodiment, and Multi-Emb denotes multi-embodiment. For fine-tuning, MT denotes multi-task training and MI denotes tasks with diverse multi-instructions. Category denotes the main capability we are trying to quantify.
Environment Category Pretraining Fine-tuning
Dataset # Trajs Dataset # Trajs
LangTable In-Domain Sim (All 5 tasks) 181k 5 Tasks (MT, MI) 1k
Cross-Task Sim (All 5 tasks) 181k 1 Task (MI) 7k
Cross-Env Real (All 5 tasks) 442k 5 tasks (MT, MI) 1k
SIMPLER In-Domain Bridgev2 60k 4 Tasks (MT) 100
Cross-Emb Something v2 200k 4 Tasks (MT) 100
Real-World Cross-Emb Bridgev2 60k 3 tasks (MI) 450
Multi-Emb Open-X 970k 3 tasks (MI) 450
Cross-Emb Something v2 200k 3 tasks (MI) 450
(a) Language Table
(b) Simpler
(c) Real
Figure 9: Experimental Setups. (a) shows an example from the 440k real-world trajectories (top) and the 181k simulation trajectories (bottom) from the Language Table Benchmark. (b) shows the 4 different evaluation tasks we use with the SIMPLER environment. (c) shows the three different tasks that we perform in the real-world.
@@
Figure 9 (C) shows examples of the real-world tabletop manipulation experimental setup. For the teleoperation, we use the polymetis robotic stack [^7] https://github.com/facebookresearch/polymetis to collect 150 trajectories for each of the tasks. All of the tasks require multi-instruction following capabilities since there are 3 objects in the scene and the model has to condition on the task description to infer which object to interact with. Figure 10 shows samples of each task. For each task, we aim to quantify 3 distinct capabilities:
(1) We test the ability to infer the correct object from the task description between an unseen combination of seen objects during fine-tuninig, (2) We test the ability to infer the correct object from totally unseen objects during fine-tuning that may or may have not been observed during pretraining. Specifically, the knocking tasks was conducted with real-world objects that were highly unlikely to have been in any of the pertaining datasets. (3) We test the ability to infer the correct object (among seen objects, unseen combinations) from a totally unseen instruction that requires semantic reasoning (e.g. Pick up a spicy object). For each evaluation criteria, 6 rollouts are performed for each models, resulting in a total of 18 rollouts for each task category. Since there are three tasks, each model is evaluated with 54 rollouts in the real-world. We provide the full list of all of the seen and unseen objects used for each rollout in Table 13, 14, 15, and the total average success rates in Table 16.
@@ Appendix D Experimental Result Analysis
We further analyze the real-world robot results shown in Figures 3 and 4(b), focusing on how the task distribution in pretraining data impacts downstream performance. Table 4 presents the number of trajectories corresponding to each evaluation task (Knocking, Covering, and Pick & Place) across pretraining datasets (Bridgev2, Open-X, and Something Something V2 (Sthv2)), determined through lexical matching. We expect future work to use other methods of analyzing the relationship between pertaining and fine-tuning task distributions that capture semantics of the task rather than simple lexical matching. We perform this analysis to get a sense of how the task distribution in the pretraining data affects downstream task performance.
@@
Appendix E Detailed Latent Action Analysis
Figure 12: Latent Action Analysis in Language Table. We condition the current observation $x_{1}$ and quantized latent action to the decoder of the latent action quantization model. We observe that each latent action can be mapped into a semantic action. For example, latent action 0 corresponds to moving a bit left and forward and corresponds to moving a bit left and back.
We provide further qualitative analysis of LAPA. First, we analyze latent actions learned from Language Table with vocabulary size of 8 and sequence length of 1. In Figure 12, we show that each latent action corresponds to a semantic action (0: Move left and forward, 1: Move left and back, 2: Move right and back, 3: Move right slightly, 4: Move right, 5: Move back, 6: Do not move, 7: Move forward). We observe that increasing the latent action vocabulary size leads to capturing a more fine-grained information.
Figure 13: Correlation of latent action with ground-truth actions When we map latent actions to ground-truth 2 DOF actions of Language Table, we observe that latent actions are well-clustered in the actual 2D action space.
We analyze the relationship between latent actions with ground-truth 2 DOF actions by mapping each instance into latent action space. As shown in Figure 13, we observe that latent actions are well-clustered in the actual 2D action space, indicating that latent actions are meaningful representations that are highly related to actual continuous actions.
We further analyze the latent actions learned from human manipulation videos using the Something-Something V2 dataset. As illustrated in Figure 14, these latent actions capture not only hand movements but also camera movements. Since the camera viewpoint varies throughout the videos in the Something-Something V2 dataset due to the videos being egocentric, our latent action quantization model also learns to represent camera movements. For instance, latent actions [3,5,2,7] and [5,6,7,6] correspond to slight downward camera movement, [4,0,0,4] and [2,3,6,6] indicate rightward movement, and [4,2,0,0] and [5,7,0,5] represent subtle upward camera shifts.
Appendix F Additional Ablation Results
We first analyze the effect of window size $H$ for latent action quantization process. For all robot manipulation videos, we have determined the window size depending on the fps of the video so that the next frame models 0.6 seconds ahead from the current frame. For human manipulation videos, we have set the next frame to be 2.4 seconds ahead since we qualitatively observed that many of the frames of the human videos contain much less dynamic actions compared to robot videos. (However, we think that filtering these frames could make the window size the same as robot videos, which we leave as future work.). We have added an ablation experiment on the window size for robot videos (Bridgev2) by evaluating on SIMPLER in Figure 15(a). Note that the default is $H=3$ because Bridgev2 is collected with 5hz. The results show that LAPA is quite robust to different window sizes. However, if the window size is extremely large, performance degradation is observed. This is expected since our quantization model is relatively small (300M parameters), it faces difficulties modeling latent information when the visual deltas are significant.
We also analyze the data scaling in terms of fine-tuning data by comparing with Scratch. We evaluate on SIMPLER. As shown in Figure 15(b), LAPA (Bridge) consistently outperforms Scratch even when the fine-tuning data instances are small while the absolute performance increases with larger fine-tuning data.
For data scaling, we also analyze the data scaling of human videos. We compare LAPA trained from 10% of Sthv2 human video dataset with LAPA trained from the whole Sthv2 human video dataset. Results in Table 12 show that scaling the human video datasets boosts the performance for SIMPLER benchmark not only for the final success for all subtasks. We leave exploring scaling law for human videos more extensively or future work, since showing scaling law requires intensive computational resources to do different ablations of model size, data size, and computational resources.
(a) Window Size
(b) Fine-tuning Data Scale
Figure 15: Additional Ablation Results of LAPA. We further analyze the performance of LAPA by varying the window size for latent action quantization and fine-tuning data scale. We report the average success rate in SIMPLER.
@@ Appendix F Additional Ablation Results
Finally, we vary the latent action length and vocabulary size in Language Table, extending the result of Figure 5 which was analyzed in Bridgev2 data. As shown in Figure 16, increasing the sequence and vocab size increases the performance. However, unlike SIMPLER, we observe that the increasing the latent action vocab size is much more effective compared to increasing the latent action sequence length in terms of absolute performance. This implies that for environments that are visually simple, increasing the latent action vocabulary might be more effective compared to sequence length.
@@
Appendix G Detailed Experimental Results
G.1 Language Table
We provide the detailed results of the experiments performed on the Language Table benchmark in Table 5, 6, 7, 8, 9, 10. For all of the tables in the appendix, we bold the best result among the comparisons and underline the second best. Each value denotes the success rate (%). 50 evaluation rollouts are performed for each task category, resulting in 250 total evaluation rollouts per model for each table.
@@ G.1 Language Table
G.2 SIMPLER
We provide results of various models evaluated on SIMPLER environment. Table 11 shows the setting where baseline models are pretrained on Bridgev2 and then finetuned on SIMPLER rollouts (100 videos). The results show detailed results for each task (stack green to yellow block, put carrot on plate, put spoon on otowel, put eggplant in basket) and subtasks (grasping and moving). As shown in Table 11, UniPi significantly underperforms all other baselines on the SIMPLER Environment. We observe that, although the generated plans from the diffusion models are quite accurate, the IDM lacks the capability to predict 7 DOF continuous actions accurately when given only 100 action-labeled trajectories. Specifically, we observe that UniPi often fails to grasp the object within the maximum step limit. This implies the effectivness of using VLAs in scenarios with insufficient action-labeled data. Similar to the results of Section 4.3, LAPA outperforms baseline models that pretrain without using ground-truth action labels (UniPi and Vpt) and closes the performance gap with ActionVLA, which is pretrained on all of the 60K action-labeled trajectories from the Bridgev2 dataset. This highlights the effectiveness of LAPA, even when the complexity of the action space increases. We also evaluate the performance of OpenVLA fine-tuned on SIMPLER trajecotries for reference. The performance of OpenVLA (36.4) is similar to Scratch. The bad performance of OpenVLA on SIMPLER is a well known issue which is due to OpenVLA not being robust to real-to-sim transfer for SIMPLER.
@@ G.2 SIMPLER
Table 11: SIMPLER results of Bridgev2 Pretraining. Success, Grasping, and Moving Rates (%) in SIMPLER environment. We pretrain UniPi, Vpt, and LAPA on Bridgev2 dataset without using ground-truth action labels and ActionVLA on Bridgev2 using action labels. We also add the result of OpenVLA finetuned on SIMPLER trajectories for reference. The main 4 tasks are: stack green to yellow block, put carrot on plate, put spoon on towel, and put eggplant in basket. Best is bolded and second best is underlined.
Success Rate Scratch UniPi Vpt LAPA ActionVLA OpenVLA
Stack G2Y 29.2 2.7 45.8 54.2 75.0 41.6
Carrot2Plate 29.2 2.7 37.5 45.8 58.0 50.0
Spoon2Towel 50.0 0.0 70.8 70.8 70.8 37.5
Eggplant2Bask 29.2 0.0 50.0 58.3 50.0 16.7
AVG 34.4 1.3 51.0 57.3 63.5 36.4
@@
Grasp Green Block 66.6 20.8 62.5 62.5 87.5 50.0
Grasp Carrot 45.8 33.2 54.1 58.3 75.0 66.6
Grasp Spoon 70.8 22.2 79.2 83.3 83.3 45.8
Grasp Eggplant 62.5 16.0 70.8 83.3 75.0 37.5
AVG 61.4 23.1 66.7 71.9 80.2 50.0
@@
Move Green Block 58.3 29.1 58.3 66.6 91.6 70.8
Move Carrot 45.8 48.6 66.6 70.8 91.6 75.0
Move Spoon 70.8 34.6 79.2 83.3 79.2 75.0
Move Eggplant 87.5 58.0 70.8 87.5 91.6 50.0
AVG 65.6 42.6 68.7 77.1 88.5 67.7
@@
Success Rate VPT UniPi LAPA LAPA (10%)
StackG2Y 50.0 0.0 50.0 45.8
Carrot2Plate 29.1 1.3 50.0 41.6
Spoon2Towel 37.5 1.3 50.0 66.6
Eggplant2Bask 66.6 0.0 58.3 45.8
AVG 45.8 0.7 52.1 50.0
@@
Grasp Green Block 66.6 2.7 58.3 50.0
Grasp Carrot 45.8 31.7 62.5 41.6
Grasp Spoon 70.8 21.7 75.0 70.8
Grasp Eggplant 91.6 6.8 70.8 62.5
AVG 68.7 15.7 66.7 56.2
@@
Move Green Block 62.5 2.7 62.5 50.0
Move Carrot 58.3 37.5 70.8 58.3
Move Spoon 54.1 18.1 75.0 79.2
Move Eggplant 91.6 50.3 83.3 62.5
AVG 66.6 27.1 72.9 62.5
G.3 Real-world
@@
OpenVLA LAPA OpenVLA LAPA ActionVLA Scratch LAPA
@@
Seen Objects, Unseen Object Combinations
@@
Seen Objects, Unseen Instructions
@@
OpenVLA LAPA OpenVLA LAPA ActionVLA Scratch LAPA
@@
Seen Objects, Unseen Object Combinations
@@
Seen Objects, Unseen Instructions
@@
OpenVLA LAPA OpenVLA LAPA ActionVLA Scratch LAPA
@@
Seen Objects, Unseen Object Combinations
@@
Seen Objects, Unseen Instructions
@@
Table 16: Summary of Total Success Rates (%)
OpenVLA LAPA OpenVLA LAPA ActionVLA Scratch LAPA